OpenStudy (anonymous):

A function is shown below: f(x) = (0.93)^x What does the function represent? Exponential growth of 7% Exponential decay of 7% Exponential growth of 93% Exponential decay of 93%

OpenStudy (freckles):

solve 1+x=.93 if x is positive then you have growth if x is negative then you have decay

OpenStudy (freckles):

have you solve 1+x=.93 for x yet?

OpenStudy (freckles):

and i probably should had used a different variable since you are already have another variable up there called x

OpenStudy (freckles):

solve 1+b=.93 for b